OpenStudy (anonymous):

3 + 6z = 13 + 6z What does z equal?

OpenStudy (anonymous):

3+6z=13+6z 6z=13+6z-3 6z=10+6z z=(10+6z)/z

OpenStudy (anonymous):

thank you @Albertoimus

OpenStudy (anonymous):

oops sorry, the last one should be z=(10+6z)/6 . sorry, other than that you should be good.

OpenStudy (turingtest):

subtract 6z from both sides

OpenStudy (anonymous):

no problem @Albertoimus thanks any way

OpenStudy (anonymous):

i already did that @TuringTest

OpenStudy (turingtest):

and what did you get? the answer provided by Albertoimus is wrong btw

OpenStudy (anonymous):

z=-10

OpenStudy (turingtest):

no that is incorrect 3 + 6z = 13 + 6z - 6z - 6z what are you left with?

OpenStudy (anonymous):

why would you subtract -6z? s/he'll end up with 3=13, which is incorrect. s/he's trying to solve for z.

OpenStudy (anonymous):

3=13+1z @TuringTest

OpenStudy (turingtest):

@rejectz48 no, what albertoinus has is the correct result of subtracting 6z from both sides @Albertoimus this does tell you the answer; namely that there is no answer that is, no real solutions since 3=13 is never true.

OpenStudy (anonymous):

so there is no solution @TuringTest

OpenStudy (anonymous):

the solution should be z=(10+6z)/6. however, if you don't have a number to replace z, then there's no solution.

OpenStudy (turingtest):

if you follow through what you have and try to isolate z you get z=(10+6z)/6 z=5/3+z 0=5/3 again you see that by trying to solve for z you get a contradiction

OpenStudy (anonymous):

ok.... @TuringTest @Albertoimus

OpenStudy (anonymous):

all in all, no solution.

OpenStudy (anonymous):

Possible solutions... x is unknown, a and b are constants x=a, one solution (a) x=x or a=a, infinite solutions a=b, no solutions